A pair of dice are rolled 1,000 times with the following frequencies of outcomes Sum 2 3 4 5 6 7 8 9 10 11 12 Frequency 10 30PART A AND B PLEASE
Answer #1

a) P(sum is less than 5 or greater than 9) = P(2) + P(3) + P(4) + P(10) + P(11) + P(12) = (10 + 30 + 50 + 120 + 80 + 70) / 1000 = 360/1000 = 0.36

Odds = 360 / (1000 - 360) = 360 / 640 = 9/16

b) P(sum is even or exactly divisible by 5) = P(2) + P(4) + P(5) + P(6) + P(8) + P(10) + P(12)

                                                                   = (10 + 50 + 70 + 110 + 170 + 120 + 70) / 1000

                                                                   = 600/1000

                                                                   = 0.6

Odds = 600 / (1000 - 600) = 600/400 = 3/2